# Towards Data-Efficient Pretraining for Atomic Property Prediction

### v1

@@ Abstract
This paper challenges the recent paradigm in atomic property prediction that links progress to growing dataset sizes and computational resources. We show that pretraining on a carefully selected, task-relevant dataset can match or even surpass large-scale pretraining, while using as little as 1/24th of the computational cost. We introduce the Chemical Similarity Index (CSI), a novel metric inspired by computer vision’s Fréchet Inception Distance, for molecular graphs which quantifies the alignment between upstream pretraining datasets and downstream tasks. By selecting the most relevant dataset with minimal CSI distance, we show that models pretrained on a smaller, focused dataset consistently outperform those pretrained on massive, mixed datasets such as JMP, even when those larger datasets include the relevant dataset. Counterintuitively, we also find that indiscriminately adding more data can degrade model performance when the additional data poorly aligns with the task at hand. Our findings highlight that quality often outperforms quantity in pretraining for atomic property prediction. Code: github.com/Yasir-Ghunaim/efficient-atom
Machine Learning, ICML
@@ 1 Introduction
Machine learning is transforming molecular modeling, driving advancements in accurate predictions and simulations of molecular behavior (Chanussot et al., 2021; Tran et al., 2023; Liao et al., 2023). These breakthroughs directly impact the acceleration of progress in crucial fields such as drug discovery (Huang et al., 2021) and global climate change mitigation (Sriram et al., 2024). The improvements in this field have been primarily attributed to innovations in model architectures (Liao et al., 2023; Gasteiger et al., 2021; Passaro & Zitnick, 2023) and the growing availability of large-scale molecular datasets. In recent years, the sizes of molecular datasets have increased dramatically - from tens of thousands of examples (Christensen & Von Lilienfeld, 2020; Chmiela et al., 2023; Wu et al., 2018) to hundreds of millions (Chanussot et al., 2021; Tran et al., 2023). This rapid growth in scale has also caused a surge in the computational resources required for pretraining, increasing from a few days on a single GPU to over a thousand GPU-days (Shoghi et al., 2023; Liao et al., 2023). This trend begs the question:
@@
Is scaling data and resources the only path forward in atomic property prediction, or can intelligent data selection achieve similar performance more efficiently?
Figure 1: Pretraining on a High-Quality, Task-Relevant Dataset. Pretraining on a carefully selected high-quality dataset achieves comparable or superior mean absolute error (MAE) across tasks while reducing computational cost by a factor of 24 compared to JMP-S, which is pretrained on all upstream datasets. Lower MAE indicates better performance.
While data selection strategies for pretraining have been explored in fields like natural language processing (Penedo et al., 2024) and computer vision (Hammoud et al., 2024; Li et al., 2023), this area remains largely underexplored in atomic property prediction, where unique challenges arise. In his paper, we challenge the prevailing assumption that ”bigger is better” by exploring whether a smaller, strategically selected dataset can lead to comparable or even superior performance while substantially reducing computational demands. We introduce a pretraining paradigm that shifts the focus from data and compute scaling to selecting the most relevant upstream dataset for improved downstream performance.
Through a simple baseline, our experiments reveal two key insights:
[…]
(2) Quality Outperforms Quantity: Expanding the pretraining dataset by incorporating additional data from less relevant sources can negatively impact downstream performance rather than enhance it.
explore the potential of dataset selection for pretraining in atomic property prediction, we introduce the Chemical Similarity Index (CSI), a simple metric inspired by the Fréchet Inception Distance (FID) from computer vision. CSI measures the alignment between an upstream dataset and a downstream task, enabling the selection of chemically relevant pretraining data. By focusing on these highly relevant datasets, we significantly reduce computational costs while maintaining competitive performance and, in many cases, achieving improvements. While large-scale datasets like OC20 (Chanussot et al., 2021; Tran et al., 2023) and mixed datasets like JMP (Shoghi et al., 2023) are popular choices for pretraining in molecular domains (Kolluru et al., 2022; Shoghi et al., 2023), our findings challenge their universal utility. Surprisingly, pretraining on a single, carefully selected dataset guided by CSI often outperforms models trained on mixtures, even when those include the most relevant dataset.
The contributions of this paper are threefold: (1) We introduce a novel framework for computationally efficient pretraining of molecular machine learning models, demonstrating that strategic data selection can match or outperform models trained on much larger datasets. (2) We propose the Chemical Similarity Index (CSI), a metric for assessing the similarity between upstream and downstream molecular datasets, enabling effective dataset selection. (3) We provide an extensive empirical evaluation demonstrating the effectiveness of our approach, offering a practical and efficient alternative to the current trend of ever-increasing data and computational costs in molecular machine learning.
@@ 2 Related Work
Pretraining for Atomic Property Prediction. Inspired by the success of pretraining in computer vision and natural language processing, pretraining for atomic property prediction has gained significant attention in recent years. Most approaches in molecular machine learning focus on self-supervised learning (Liu et al., 2021; Jiao et al., 2023; Chen et al., 2021; Kolluru et al., 2022; Zhou et al., 2022a; Ji et al., 2024), as generating labels for molecular datasets is computationally expensive. In contrast, fewer studies explore the effectiveness of supervised transfer learning (Smith et al., 2019, 2018; Kolluru et al., 2022). However, in both self-supervised and supervised settings, the focus has primarily been on improving feature representation, often overlooking the impact of pretraining dataset relevance on downstream performance. Recently, JMP (Shoghi et al., 2023) introduced multi-domain pretraining, enabling joint pretraining on various upstream sources concurrently. While effective, JMP pretraining requires enormous computational resources to reproduce and does not reveal how each upstream source impacts downstream performance. Our work addresses this gap by systematically studying the relationship between upstream pretraining datasets and downstream performance, enabling researchers to develop effective pretraining models even with limited computational resources.
@@
Data Selection. Efficient training through data selection has been explored via two primary approaches: subset selection and dataset distillation. Subset selection aims to identify a representative subset of the training data that matches or even outperforms training on the full dataset. Several methods have been proposed for vision and NLP tasks (Attendu & Corbeil, 2023; Killamsetty et al., 2021a, b; Kaushal et al., 2019; Bairi et al., 2015; Lapedriza et al., 2013). Dataset distillation, introduced by Wang et al. (Wang et al., 2018), focuses on generating a smaller, synthetic subset of the dataset that preserves performance while reducing training time and storage requirements. Subsequent work has explored techniques such as meta-learning (Zhou et al., 2022b; Nguyen et al., 2021a, b), gradient matching (Zhao et al., 2021), and distribution matching (Zhao & Bilen, 2023). While most research in distillation has focused on vision tasks, a few studies have extended it to graph data (Jin et al., 2022b; Liu et al., 2022; Jin et al., 2022a), though primarily targeting knowledge and social graphs rather than molecular graphs.
@@ 3 Formulation and Setup
In this section, we present our problem setup, notion of a computational budget, and the a formulation of dataset similarity. We then detail how we adapt the Fréchet Inception Distance (FID) to the molecular domain, yielding the Chemical Similarity Index (CSI). Our setup is illustrated in Figure 2. Throughout this work, we use the term ’molecular’ broadly to encompass both molecular and materials domains, as well as their respective datasets.
@@ 3.1 Formal Setting
Let $\{\mathcal{D}_{u}^{(1)},\mathcal{D}_{u}^{(2)},\ldots,\mathcal{D}_{u}^{(K)}\}$ denote a collection of $K$ upstream (pretraining) datasets, each containing molecular structures paired with relevant atomic properties (e.g., energies and forces). In the typical paradigm, the user aggregates all upstream datasets into one dataset for pretraining:
$\mathcal{D}_{u}=\bigcup_{i=1}^{K}\mathcal{D}_{u}^{(i)}.$
@@
Often, $\mathcal{L}_{\text{energy}}$ is computed using the Mean Absolute Error (MAE), while $\mathcal{L}_{\text{forces}}$ uses the mean per-atom Euclidean (L2) distance. Coefficients $\alpha$ and $\beta$ weight the importance of energy and force tasks, respectively, with $\beta>\alpha$ often chosen to prioritize accurate force predictions in atomistic modeling. Usually, pretraining is performed on the joint upstream dataset $\mathcal{D}_{u}$ or a subsampled version.
@@
In this paper, we consider two additional needed definitions for this setting: (1) computational budget and (2) similarity metrics.
@@
Hence, the computational budget $\mathcal{C}$ represents the total number of samples processed over training. It naturally splits into two factors: the dataset size ($\mathcal{N}$) and the number of passes through it ($\mathcal{E}$). The choice of $\mathcal{C}$ depends on the available computing resources. In our experiments, we explore multiple pretraining budget settings to examine how trading off $\mathcal{N}$ against $\mathcal{E}$ under a fixed $\mathcal{C}$ affects both pretraining and downstream performance.
@@
that quantifies their relevance or “similarity.” In principle, a smaller value $\delta(\mathcal{D}_{u},\mathcal{D}_{d})$ indicates a greater overlap or relevance. Thus, among multiple candidate upstream datasets $\{\mathcal{D}_{u}^{(1)},\ldots,\mathcal{D}_{u}^{(K)}\}$, the one that minimizes
@@
should provide the most effective pretraining for $\mathcal{D}_{d}$. In this paper, we empirically test this assumption, examining whether lower $\delta$-values indeed correlate with improved downstream performance. This motivates using $\delta$ as a principled metric to select or combine upstream datasets under a fixed computational budget.
@@ 3.2 The Chemical Similarity Index (CSI)
Our proposed Chemical Similarity Index (CSI) draws its inspiration from the well-known Fréchet Inception Distance (FID)  (Heusel et al., 2017). Recall that FID is commonly used in computer vision to compare two sets of images via their feature distributions. Specifically, if one extracts features (e.g., from an Inception network) for datasets $X$ and $Y$ and denotes their empirical means and covariances by $\mu_{X},\Sigma_{X}$ and $\mu_{Y},\Sigma_{Y}$, then
@@
Figure 3: Alignment Between Upstream and Downstream Using CSI. We assess how well the extracted representations from each upstream dataset align with downstream tasks using our CSI metric, where lower values indicate stronger alignment. ANI-1x demonstrates the closest feature alignment with downstream tasks, whereas OC20 and OC22 show the weakest alignment.
Design Considerations for CSI.
Adapting FID to the molecular domain involves several key design choices regarding how to best modify the metric:
1. Feature Type: The original FID metric for images uses features from the last layer before the head, relying on a single feature type. In contrast, molecular graph-based models often use two types of features: node and edge features. Models like GemNet-OC (Gasteiger et al., 2022) retain distinct node and edge features before the head, while transformer-based models like EquiformerV2 (Liao et al., 2023) integrate edge information into unified node embeddings. To ensure compatibility across backbones, we focus on node embeddings at the last backbone layer, as edge information is either explicitly maintained (e.g., GemNet-OC) or indirectly encoded into node representations (e.g., EquiformerV2).
2. Aggregation Strategy: FID in computer vision operates on fixed-size feature vectors, while molecular graphs have variable-sized embeddings. To adapt FID, we consider two strategies: flattening, which concatenates all node features into a single vector, and mean pooling, which averages node features for a graph-level representation. While mean pooling aligns with the original FID approach, it may oversmooth molecular features. In models like EquiformerV2 (Liao et al., 2023) and GemNet-OC (Gasteiger et al., 2022), backbone node features are designed for further processing in the model’s head before pooling. Since we extract features directly from the backbone, mean pooling at this stage risks losing feature diversity. Thus, we choose flattening to preserve the full feature distribution. CSI values using mean pooling are shown in the appendix.
3. Sampling in Long-Tail Distributions: To make feature extraction computationally feasible, we subsample a representative subset from each upstream and downstream dataset. However, molecular datasets often exhibit long-tail distributions, where common structures are overrepresented and rare ones underrepresented. These imbalances can skew mean and covariance estimates for our metric. For example, specific molecules may appear much more frequently than others, and in catalytic datasets, certain slab–adsorbate configurations can dominate (Chanussot et al., 2021). To address this, we use a class-balancing sampling scheme to ensure a balanced representation of structures. Plots for long-tail distribution are in the appendix.
Overall, our CSI metric leverages node embeddings with flattening for aggregation and uses class-balanced sampling to address distribution imbalances. We subsample 10k instances from each upstream and downstream dataset for feature extraction to maintain computational feasibility. To ensure our metric remains independent of the baselines in this study, we extract features for our metric using EquiformerV2 (Liao et al., 2023) pretrained on OC20 (Chanussot et al., 2021). For completeness, CSI values computed using the baseline JMP (Shoghi et al., 2023) are included in the appendix.
@@
In Figure 3, we present the CSI values for pairs of upstream and downstream tasks related to energy and force predictions, with additional details about the datasets and targets provided in Section 4. ANI-1x (Smith et al., 2020) consistently achieves the closest alignment across all downstream tasks, reflecting its design goal of maximizing chemical diversity. Transition-1x (Schreiner et al., 2022), which focuses on transition states, shows as the second most aligned dataset, suggesting that its emphasis on high-energy transition states leads to partial overlap with downstream distributions. In contrast, the catalysis datasets, OC20 (Chanussot et al., 2021) and OC22 (Tran et al., 2023), exhibit the weakest alignment. While OC20 and OC22 are often favored for pretraining (Shoghi et al., 2023; Kolluru et al., 2022) due to their scale and chemical diversity, our metric suggests they may not align well with the considered downstream tasks. Next, we examine whether these alignment values correlate with downstream performance.

### v2

# On the Importance of Pretraining Data Alignment for Atomic Property Prediction

@@ Abstract
This paper challenges the recent paradigm in atomic property prediction that links progress to growing dataset sizes and computational resources. We show that pretraining on a carefully selected task-aligned dataset can match or even surpass large-scale joint pretraining while using only 1/24th of the pretraining budget. We introduce the Chemical Similarity Index (CSI), a simple metric for molecular graphs inspired by the Fréchet Inception Distance in computer vision, which quantifies the alignment between upstream pretraining datasets and downstream tasks. By selecting the most aligned dataset with minimal CSI distance, we show that models pretrained on a smaller, focused dataset consistently achieve better performance on downstream tasks than those pretrained on massive, mixed datasets such as JMP. This holds even when the mixed dataset includes the upstream dataset most aligned with the downstream task. Counterintuitively, we also find that indiscriminately adding more data can degrade model performance when the additional data is poorly aligned with the target task. Our findings highlight that quality often outperforms quantity in pretraining for atomic property prediction. The code is publicly available at: github.com/Yasir-Ghunaim/efficient-atom.
@@ 1 Introduction
Machine learning is transforming molecular modeling, driving advancements in accurate predictions and simulations of molecular behavior (Chanussot et al., 2021; Tran et al., 2023; Liao et al., 2023). These breakthroughs directly impact the acceleration of progress in crucial fields such as drug discovery (Huang et al., 2021) and global climate change mitigation (Sriram et al., 2024). The improvements in this field have been primarily attributed to innovations in model architectures (Liao et al., 2023; Gasteiger et al., 2021; Passaro and Zitnick, 2023) and the growing availability of large-scale molecular datasets. In recent years, the sizes of molecular datasets have increased dramatically, from tens of thousands of examples (Christensen and Von Lilienfeld, 2020; Chmiela et al., 2023; Wu et al., 2018) to hundreds of millions (Chanussot et al., 2021; Tran et al., 2023). This rapid growth in scale has also caused a surge in the computational resources required for pretraining, increasing from a few days on a single GPU to over a thousand GPU-days (Shoghi et al., 2023; Liao et al., 2023). This trend raises the question:
Is scaling data and resources the only path forward for pretraining in atomic property prediction, or can intelligent data selection achieve similar performance more efficiently?
Figure 1: Pretraining on a High-Quality, Task-Aligned Dataset. Pretraining on a carefully selected high-quality dataset achieves comparable or superior mean absolute error (MAE) across tasks while reducing pretraining budget by a factor of 24 compared to JMP-S, which is pretrained on all upstream datasets. Lower MAE indicates better performance.
While data selection strategies for pretraining have been explored in fields like natural language processing (Penedo et al., 2024) and computer vision (Hammoud et al., 2024; Li et al., 2023), this area remains largely underexplored in atomic property prediction, where unique challenges arise. This gap is especially critical for 3D molecular and material pretraining, which requires far more computation (Liao et al., 2023) than methods based on 2D representations (Rong et al., 2020). In this paper, we challenge the prevailing assumption that "bigger is better" by exploring whether a smaller, strategically selected dataset can lead to comparable or even superior performance while substantially reducing computational demands. We introduce a pretraining paradigm for 3D molecular and material systems that shifts the focus from data and compute scaling to selecting the most relevant upstream dataset for improved downstream performance.
@@
To explore the potential of dataset selection for pretraining in atomic property prediction, we introduce the Chemical Similarity Index (CSI), a simple metric inspired by the Fréchet Inception Distance (FID) from computer vision. CSI measures the alignment between an upstream dataset and a downstream task, enabling the selection of chemically relevant pretraining data. By focusing on these highly relevant datasets, we significantly reduce computational costs while maintaining competitive performance and, in many cases, achieving improvements. While large-scale datasets like OC20 (Chanussot et al., 2021; Tran et al., 2023) and mixed datasets like JMP (Shoghi et al., 2023) are popular choices for pretraining in molecular domains (Kolluru et al., 2022; Shoghi et al., 2023), our findings challenge their universal utility. Surprisingly, pretraining on a single, carefully selected dataset guided by CSI often outperforms models trained on mixtures, even when those include the most relevant dataset.
The contributions of this paper are threefold: (1) We introduce a novel framework for computationally efficient pretraining of molecular machine learning models, demonstrating that strategic data selection can match or outperform models trained on much larger datasets. (2) We propose the Chemical Similarity Index (CSI), a metric for assessing the alignment between upstream and downstream molecular datasets, enabling effective dataset selection. (3) We provide an extensive empirical evaluation demonstrating the effectiveness of our approach, offering a practical and efficient alternative to the current trend of ever-increasing data and computational costs in molecular machine learning.
@@ 2 Related Work
Pretraining for Atomic Property Prediction. Inspired by the success of pretraining in computer vision and natural language processing, pretraining for atomic property prediction has gained significant attention in recent years. Most approaches in molecular machine learning have focused on self-supervised pretraining (Rong et al., 2020; Liu et al., 2021; Jiao et al., 2023; Chen et al., 2021; Zhou et al., 2022a; Ji et al., 2024), while fewer studies have explored the effectiveness of supervised pretraining (Smith et al., 2019; 2018; Kolluru et al., 2022). Early self-supervised methods such as GROVER (Rong et al., 2020) target 2D molecular graphs. More recent approaches, including GraphMVP (Liu et al., 2021), Uni-Mol (Zhou et al., 2022a) and Uni-Mol2 (Ji et al., 2024), extend to 3D structures and often use denoising objectives that are best suited for equilibrium geometries (Liao et al., 2024). This limits their applicability to large-scale non-equilibrium datasets common in molecular and materials modeling. In contrast, recent supervised pretraining frameworks such as Joint Multi-domain Pre-training (JMP) (Shoghi et al., 2023) jointly pretrain on diverse large-scale labeled datasets and operate effectively on non-equilibrium data. While powerful, JMP requires substantial compute due to the massive data volume and does not reveal how each pretraining dataset influences downstream performance. Our work addresses this gap by systematically analyzing the link between pretraining dataset choice and downstream performance. Based on our analysis, we show that selecting a single task-aligned upstream dataset can match or exceed joint pretraining, enabling efficient model development under constrained computational budgets.
@@
Distribution Similarity. A variety of methods exist for quantifying the distance between probability distributions, including Kullback–Leibler (KL) divergence (Kullback and Leibler, 1951), Jensen–Shannon (JS) divergence (Lin, 2002), Maximum Mean Discrepancy (MMD) (Gretton et al., 2012), and the Fréchet Inception Distance (FID) (Heusel et al., 2017). FID has become a standard metric in computer vision for comparing real and generated data distributions by evaluating differences in the means and covariances of their feature representations. In the chemical domain, the Fréchet ChemNet Distance (FCD) (Preuer et al., 2018) adapts FID for evaluating generative models that use SMILES-based molecular representations and has been integrated into benchmarking frameworks such as GuacaMol (Brown et al., 2019). While FCD demonstrates the applicability of FID-based metrics in chemistry, it is primarily applied to SMILES-based molecular data within a single chemical domain. Our proposed Chemical Similarity Index (CSI) is also derived from FID, but is designed for 3D graph-structured systems and focuses on measuring alignment across multiple domains such as molecules and materials. To our knowledge, we are the first to systematically study pretraining dataset alignment in 3D atomic property prediction.
Data Selection. Efficient training through data selection has been explored via two primary approaches: subset selection and dataset distillation. Subset selection aims to identify a representative subset of the training data that matches or even outperforms training on the full dataset. Several methods have been proposed for vision and NLP tasks (Attendu and Corbeil, 2023; Killamsetty et al., 2021a; b; Kaushal et al., 2019; Bairi et al., 2015; Lapedriza et al., 2013). Dataset distillation, introduced by Wang et al. (2018), focuses on generating a smaller, synthetic subset of the dataset that preserves performance while reducing training time and storage requirements. Subsequent work has explored techniques such as meta-learning (Zhou et al., 2022b; Nguyen et al., 2021a; b), gradient matching (Zhao et al., 2021), and distribution matching (Zhao and Bilen, 2023). While most research in distillation has focused on vision tasks, a few studies have extended it to graph data (Jin et al., 2022b; Liu et al., 2022; Jin et al., 2022a), though primarily targeting knowledge and social graphs rather than molecular graphs.
@@ 3 Formulation and Setup
In this section, we present our problem setup, notion of a computational budget, and the formulation of dataset similarity. We then detail how we adapt the Fréchet Inception Distance (FID) to the molecular domain, yielding the Chemical Similarity Index (CSI). Our setup is illustrated in Figure 2. Throughout this work, we use the term ’molecular’ broadly to encompass both molecular and materials domains, as well as their respective datasets.
@@ 3.1 Formal Setting
Let $\{\mathcal{D}_{u}^{(1)},\mathcal{D}_{u}^{(2)},\ldots,\mathcal{D}_{u}^{(K)}\}$ denote a collection of $K$ upstream (pretraining) datasets, each containing molecular structures paired with relevant atomic properties (e.g., energies and forces). In the typical paradigm, upstream datasets are typically aggregated into a single pretraining set:
$\mathcal{D}_{u}=\bigcup_{i=1}^{K}\mathcal{D}_{u}^{(i)}.$ (1)
@@
We compute $\mathcal{L}_{\text{energy}}$ using the Mean Absolute Error (MAE) and $\mathcal{L}_{\text{forces}}$ using the mean per-atom Euclidean (L2) distance. Coefficients $\alpha$ and $\beta$ weight the importance of energy and force tasks, respectively. Following the JMP paper (Shoghi et al., 2023), we set $\beta>\alpha$ to prioritize accurate force predictions in atomistic modeling. Pretraining can be performed on either the joint upstream dataset $\mathcal{D}_{u}$, similar to JMP (Shoghi et al., 2023), or on an individual upstream dataset $\mathcal{D}_{u}^{(i)}$, as in our selective setting.
@@
In this paper, we introduce two additional definitions for this setting: (1) computational budget and (2) dataset similarity.
@@
Hence, the computational budget $\mathcal{C}$ represents the total number of samples processed over training. It naturally splits into two factors: the dataset size ($\mathcal{N}$) and the number of passes through it ($\mathcal{E}$). The choice of $\mathcal{C}$ depends on the available computing resources. In our main experiments, we fix $\mathcal{C}$, $\mathcal{N}$, and $\mathcal{E}$ to ensure a fair comparison across different upstream datasets. We also include experiments in which $\mathcal{N}$ (and thus $\mathcal{C}$) varies, in order to analyze the impact of dataset size and total compute on downstream performance.
@@
that quantifies their alignment or “similarity.” In principle, a lower value of $\delta(\mathcal{D}_{u},\mathcal{D}_{d})$ reflects a higher degree of alignment between the upstream and downstream distributions. Thus, among multiple candidate upstream datasets $\{\mathcal{D}_{u}^{(1)},\ldots,\mathcal{D}_{u}^{(K)}\}$, the one that minimizes
@@
should provide the most effective pretraining for $\mathcal{D}_{d}$. In this paper, we empirically test this assumption, examining whether lower $\delta$-values indeed correlate with improved downstream performance. Motivated by this, we use $\delta$ as a principled metric to guide dataset selection for Eq. (1). Instead of aggregating all upstream datasets, we modify the pretraining setup to use only the single dataset $\mathcal{D}_{u}^{(i)}$ that best aligns with the downstream task under a fixed computational budget.
@@ 3.2 The Chemical Similarity Index (CSI)
Our proposed Chemical Similarity Index (CSI) draws its inspiration from the well-known Fréchet Inception Distance (FID(Heusel et al., 2017). FID is commonly used in computer vision to compare two sets of images via their feature distributions. Specifically, if one extracts features (e.g., from an Inception network) for datasets $X$ and $Y$ and denotes their empirical means and covariances by $\mu_{X},\Sigma_{X}$ and $\mu_{Y},\Sigma_{Y}$, then
@@
To adapt FID for graph-structured molecular data, we compute the CSI metric using node embeddings as features and apply class-balanced sampling to ensure representative coverage of molecular types in each upstream dataset. For computational feasibility, we subsample 10k instances from both the upstream and downstream datasets. To keep the metric fixed and comparable across all evaluations, we extract features using EquiformerV2 (Liao et al., 2023) pretrained on OC20 (Chanussot et al., 2021). We provide results with additional pretrained checkpoints in Appendix B, feature representation choices in Appendix C, and class-balancing procedure in Appendix D.
@@
In Figure 3, we present the CSI values for pairs of upstream and downstream tasks related to energy and force predictions, with additional details about the datasets and targets provided in Section 4. For the first four downstream datasets in Figure 3, ANI-1x (Smith et al., 2020) consistently achieves the closest alignment, reflecting its design goal of maximizing chemical diversity. Transition-1x (Schreiner et al., 2022), which focuses on transition states, ranks second suggesting that its emphasis on high-energy transition states leads to partial overlap with downstream distributions. OMat24 (Barroso-Luque et al., 2024), which contains inorganic materials, is more closely aligned with the catalysis datasets OC20 (Chanussot et al., 2021) and OC22 (Tran et al., 2023) than with other upstream datasets. While OC20 and OC22 are often favored for pretraining (Shoghi et al., 2023; Kolluru et al., 2022) due to their scale and chemical diversity, our metric suggests they may not align well with most of the considered downstream tasks. Next, we examine whether these alignment values correlate with downstream performance.
Figure 3: Alignment Between Upstream and Downstream Using CSI. We assess how well the extracted representations from each upstream dataset align with downstream tasks using our CSI metric, where lower values indicate stronger alignment. ANI-1x demonstrates the closest feature alignment with most downstream tasks. However, for OMat24, the catalysis datasets OC20 and OC22 exhibit the strongest alignment.
4 Experiments
Downstream Datasets: For downstream evaluation, we focus on in-distribution (ID) tasks involving energy or force prediction, following the definition in JMP (Shoghi et al., 2023). We discuss out-of-distribution (OOD) tasks in Section 5. To keep the experiments tractable, we evaluate on a single target per downstream dataset, since we run six baselines for each. These targets and their corresponding datasets are: Aspirin in rMD17 (Christensen and Von Lilienfeld, 2020), Ac-Ala3-NHMe in MD22 (Chmiela et al., 2023), solvated amino acids in SPICE (Eastman et al., 2023), $U_{0}$ in QM9 (Wu et al., 2018), and force prediction on the rattled-300-subsampled split of OMat24 (Barroso-Luque et al., 2024).
Baselines: We report the original performance of JMP, where "JMP-S" and "JMP-L" correspond to the small and large backbones, respectively. Additionally, we present our reproduced fine-tuning results using the official JMP checkpoints, denoted as "JMP-S*" and "JMP-L*".
For our budgeted evaluation, we present results in two categories: pretraining on a single upstream dataset and pretraining on a joint combination of all upstream datasets. For single-dataset experiments, we randomly sample $\mathcal{N}$ instances from the original upstream data. For joint pretraining, we construct the training set using two different strategies. (1) Balanced Sampling, where an equal number of samples is drawn from each of the four upstream datasets, totaling $\mathcal{N}$ samples; and (2) Temperature-Based Sampling, which preserves the dataset proportions used in the full 120M sample set of JMP (Shoghi et al., 2023).
Evaluation Setup: We pretrain the GemNet-OC-S model (Gasteiger et al., 2022) on each individual upstream dataset, as well as on joint configurations that combine all upstream datasets, following the baseline setups. For our main experiments, we set a fixed computational budget of $\mathcal{C}=10$M, achieved by training on $\mathcal{N}=2$M samples for $\mathcal{E}=5$ epochs. This represents a $24\times$ reduction in pretraining budget compared to that used in JMP (Shoghi et al., 2023). Our budget ensures accessibility and reproducibility, with each pretraining run completing within 1 to 2 days on an A100 GPU. Additional budget configurations are explored in later sections and the appendix. Each pretrained model is then fine-tuned separately on each downstream task.
4.1 Does CSI Correlate with Better Performance?
In Figure 3, we presented CSI values quantifying the alignment between each upstream and downstream dataset. These results raise a critical question:
Table 1 summarizes the downstream performance of models pretrained on different datasets in the in-distribution setting. Consistent with the CSI values, the upstream dataset with the lowest CSI for each downstream task also achieves the best fine-tuning performance. Specifically, pretraining on ANI-1x, despite being the smallest dataset, yields the best fine-tuning results on rMD17, MD22, SPICE and QM9, outperforming all other individual pretraining datasets as well as the joint variants. For instance, on the rMD17, SPICE and QM9 datasets, the model pretrained on ANI-1x achieves MAEs of 5.2, 5.39 and 3.1, compared to 6.7, 5.71 and 3.3 for JMP-S. This strong performance is achieved with less than 5% of the pretraining budget used by JMP-S. For OMat24, the model pretrained on OC20 achieves the best performance among the limited-budget baselines with an MAE of 87.4, comparable to JMP-S’s 84.4 despite its much larger 240M pretraining budget. This outcome is also predicted by CSI, which ranks OC20 as the most similar upstream dataset for this task.